XYZ Company recorded the following information related to their inventory accounts for 2020: January 1, 2020 December 31, 2020 Direct materials 37,000 50,000 Work in process 38,000 41,000 Finished goods 22,000 34,000 The following information was taken from XYZ Company's accounting records for 2020: Sales revenue ........................................... $630,000 Direct materials purchased .............................. ? Depreciation, factory equipment ......................... 34,000 Prime costs ............................................. 250,000 Utilities (60% for factory; 40% for office building) .... 20,000 Sales commissions ....................................... 29,000 Indirect materials ...................................... 25,000 Depreciation, office equipment .......................... 30,000 Rent, factory building .................................. 56,000 Direct labor ............................................ ? Advertising ............................................. 75,000 Production supervisor's salary .......................... 81,000 Additional information: 1. Direct labor comprised 35% of the conversion costs for 2020. 2. The actual overhead cost for 2020 was equal to the overhead applied to production. Thus there was no overhead variance for 2020. Calculate the direct materials purchased by XYZ Company in 2020.
Answer #1

Overhead applied:

Depreciation, factory equipment $34,000
Utilities (20,000*60%) 12,000
Indirect material 25,000
Rent, factory building 56,000
Production supervisor salary 81,000
Overhead applied to production $208,000

Conversion cost = Direct labor+overhead applied

Direct labor = 35% of conversion cost

Overhead applied = 65% of conversion cost

Conversion cost = $208,000/65% = $320,000

Direct labor = $320,000*35% = $112,000

Prime cost = 250,000

Prime cost = Direct material used + Direct labor

Direct material used = $250,000-112,000 = $138,000

Direct material purchased = $138,000+50,000-37,000 = $151,000